OpenStudy (anonymous):

the greater of two numbers is 10more than the smaller. their sum is 38. Find the numbers?

OpenStudy (sandra):

ok, so try and turn the first piece of information into an equation.

OpenStudy (anonymous):

how do i know which is x and y?

OpenStudy (sandra):

well, when you think about it, it's arbitrary. just pick in your head. so for this problem, let's say the greater number is X, and the smaller number is Y

OpenStudy (anonymous):

x=y=10?

OpenStudy (anonymous):

x= y +10?

OpenStudy (sandra):

exactly =)

OpenStudy (sandra):

ok , so now try and turn the second piece of information into an equation. since again, you need two equations if you have two unknowns

OpenStudy (anonymous):

y+10+y=38 2y+10=38 2y=28 ???

OpenStudy (sandra):

!! great =)

OpenStudy (sandra):

and...

OpenStudy (anonymous):

y=14

OpenStudy (sandra):

exactly =)

OpenStudy (sandra):

so what's x?

OpenStudy (anonymous):

24?

OpenStudy (sandra):

awesome job

OpenStudy (sandra):

and yes

OpenStudy (sandra):

so X, the greater number is 24, and y, the smaller number is 14

OpenStudy (anonymous):

wow! It is much easier when there is someone helping

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

I am going to work on a few more now while it is fresh in my head

OpenStudy (sandra):

no problem! good luck!

OpenStudy (sandra):

yep you got it =)

OpenStudy (anonymous):

my next problem says the greater of two numbers is 3 times the smaller. Would that meant I subtract?

OpenStudy (sandra):

nope. it basically means x = 3 * y (times)

OpenStudy (anonymous):

ok. thank you

OpenStudy (sandra):

yep np!

OpenStudy (sandra):

ok what's up

OpenStudy (sandra):

here